OpenStudy (anonymous):

is this write Estimate by rounding to the nearest whole number. 27 2/11 divided 2 5/7 =9

OpenStudy (anonymous):

\[\frac{ 27+\frac{ 2 }{ 11 } }{ 2+ \frac{ 5 }{ 7 }} ?\]